Circular No. 5827

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET WHIPPLE (1993n) J. V. Scotti, University of Arizona, reports his recovery of this comet

with the Spacewatch telescope at Kitt Peak. The comet appears essentially stellar, and the following positions are very close to the prediction on MPC 18259 (ephemeris on MPC 21961):

1993 UT  α_{2000}  δ_{2000}  m_{1}
June 25.38638  19^{h}51^{m}43.^{s}55  −8^{o}25′59′′.9
25.39787  19 51 43.11  −8 26 00.3  21.5
25.41126  19 51 42.63  −8 25 59.9  21.0
26.34800  19 51 13.58  −8 26 25.8  21.2
26.35657  19 51 13.28  −8 26 25.9  20.7
26.36824  19 51 12.87  −8 26 26.9

SUPERNOVA 1993T IN ANONYMOUS GALAXY E. Cappellaro and M. Della Valle, European Southern Observatory, La

Silla, report: “On June 28.4 UT we obtained a low-resolution spectrogram (range 400-900 nm, resolution 2 nm) of SN 1993T (cf. IAUC 5812) using the New Technology Telescope (+ EMMI). The spectrum shows several broad emission lines at 500, 535, 605, 630, and 710 nm, indicating a type-Ia supernova ∼ 40 days after maximum. The redshift of the parent galaxy, as measured from several absorption features, is z = 0.088.”

NOVA IN M31 A. V. Filippenko and T. Matheson, University of California at Berkeley,

report that a CCD spectrum (range 420-710 nm) obtained on June 28 UT with the Shane 3-m reflector at Lick Observatory confirms that the object discovered by Treffers et al. (IAUC 5814) is indeed a nova. Relatively broad, double-peaked Hα emission is detected. The nova is now much fainter than at the time of discovery.

RY SAGITTARII This R CrB-type variable (cf. IAUC 5084) has started to fade slowly for

the first time since 1990, as indicated by the following visual observations: June 1.8 UT, 5.9 (A. Jones, Nelson, New Zealand); 15.0, 6.5 (D. Overbeek, Edenvale, South Africa); 15.4, 6.9 (Jones); 17.42, 7.2 (C. Henshaw, Maunatlala, Botswana); 26.8, 7.4 (Jones); 29.1, 7.6 (Overbeek).

1993 June 30 Daniel W. E. Green